OpenStudy (anonymous):

solve 2cos^2(3x)+3cos(3x)+1=0 for x=[0,2pi)

OpenStudy (fibonaccichick666):

you have a quadratic in cosine

OpenStudy (fibonaccichick666):

if we said let u=cos3x. WHat would your eq become?

OpenStudy (anonymous):

it would be 2cos^2(u)+3cos(u)+1=0

OpenStudy (fibonaccichick666):

not quite \[\color\red{cos 3x}=u\]

OpenStudy (anonymous):

oh 2u^2+3+1=0

OpenStudy (fibonaccichick666):

well 3u but very close

OpenStudy (anonymous):

omg, im sorry.

OpenStudy (anonymous):

2u^2+3u+1=0

OpenStudy (fibonaccichick666):

it's cool , ok so now could you solve that for u?

OpenStudy (fibonaccichick666):

and don't worry, first time I saw one of these god awful things I was beyond clueless

OpenStudy (anonymous):

1/2 or -1/2 ?

OpenStudy (fibonaccichick666):

wolfram agrees! http://www.wolframalpha.com/input/?i=solve+2u%5E2%2B3u%2B1%3D0

OpenStudy (fibonaccichick666):

ok so now, you need to substitute your u value back in and solve both outcomes

OpenStudy (anonymous):

And when it equals 0 it is that one? so it would be -1/2?

OpenStudy (fibonaccichick666):

wait wheredid that come from?

OpenStudy (anonymous):

i plugged in (-1/2) and (1/2) for U and then i got 0 and 3 when plugged in

OpenStudy (fibonaccichick666):

it's just what was the substitution we had for u?

OpenStudy (anonymous):

u is cos3x

OpenStudy (anonymous):

so would it then be 2(-1/2)^2+3(-1/2)+1 ?

OpenStudy (fibonaccichick666):

uhm, just wait a sec, so u=cos3x yeah? u also equals 1/2 and -1/2

OpenStudy (fibonaccichick666):

you shouldn't get zero ever. Can you show me your work?

OpenStudy (fibonaccichick666):

ok that's why, you should solve cos3x=1/2 and cos3x=-1/2

OpenStudy (anonymous):

i think i am still doing it wrong because then it would be cos3x= 60 degrees or 120 degrees? and then do you divide that by 3 so then it would be 20 degrees or 40 degrees?

OpenStudy (fibonaccichick666):

no, ok so if you have 5x= 4 what do you do?

OpenStudy (anonymous):

its x=4/5

OpenStudy (fibonaccichick666):

ok so now you have cos(5x)=4, what do you have to do first?

OpenStudy (fibonaccichick666):

you can't just divide the 5

OpenStudy (anonymous):

cos^-1(4)=5x?

OpenStudy (anonymous):

then divide by..5?

OpenStudy (fibonaccichick666):

Heloo new person..?

OpenStudy (fibonaccichick666):

Let's wait for _lostmathstudent to figure it out

OpenStudy (anonymous):

but thats what i did

OpenStudy (anonymous):

i did the inverse and got 60 for the 1/2

OpenStudy (anonymous):

and then so for the -1/2 it could be either 120 or 240 or 300?

OpenStudy (fibonaccichick666):

by you then get 3x=60. not just 60

OpenStudy (anonymous):

so then i'd divide all of them by 3 and then get 20, 40, 80 and 100?

OpenStudy (fibonaccichick666):

wait wait wait

OpenStudy (anonymous):

okay sorry

OpenStudy (fibonaccichick666):

you are so not done with the first part, one step at a time here.

OpenStudy (fibonaccichick666):

\[cos(3x)=.5\] \[cos^{-1}[cos(3x)]=cos^{-1}(.5)\] Do you follow?

OpenStudy (anonymous):

okay i follow.

OpenStudy (fibonaccichick666):

so now, we simplify yes? so when is the cosine positive?

OpenStudy (anonymous):

in q1

OpenStudy (fibonaccichick666):

anywhere else?

OpenStudy (anonymous):

and q4

OpenStudy (fibonaccichick666):

yep ok so now, what is the value in 1 and what is the value in 4 of arccos(.5)?

OpenStudy (anonymous):

it'd be 60 and 300?

OpenStudy (fibonaccichick666):

yep, now, you have two eqs. 3x=60 and 3x=300

OpenStudy (anonymous):

so then its 20 or 100?

OpenStudy (fibonaccichick666):

both, you will have 4 answers here

OpenStudy (fibonaccichick666):

so now, you repeat for -.5

OpenStudy (anonymous):

so its 20 and 100 and then for the negative cosine it would be 120 and 240 and so then the final 4 answers would be 20,100, 40, and 80?

OpenStudy (fibonaccichick666):

sounds right

OpenStudy (anonymous):

awesome thank you !

OpenStudy (fibonaccichick666):

np, do you understand now?

OpenStudy (anonymous):

yes i do :)

OpenStudy (fibonaccichick666):

yay!